OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 7) and (–1, 3)? (1 point)

OpenStudy (strawberry17):

y2 - y1 m = ------ x2 - x1

OpenStudy (strawberry17):

So plug in the values: y2 = 3 y1 = -1 x2 = 7 x1 = 2

OpenStudy (strawberry17):

the first equation is to find the slope. So: y2 - y1 3 - (-1) 4 m = ------ = ------- = -- x2 - x1 7 - 2 5 m = 4/5 So the slope the line is 4/5.